xx TestArrays

TestArrays

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11 tghost4d 10 Scope: private01  INT

11 Description: Ghostzone size in t direction for 4d arrays

11 Range 10              01  Default: 1

11 0:*  10

11 tsize4d 10 Scope: private01  INT

11 Description: Size of 4d array in t direction

11 Range 10              01  Default: 2

11 0:*  10

11

11 xghost1d 10 Scope: private01  INT

11 Description: Ghostzone size in x direction for 1d arrays

11 Range 10              01  Default: 1

11 0:*  10

11

11 xghost2d 10 Scope: private01  INT

11 Description: Ghostzone size in x direction for 2d arrays

11 Range 10              01  Default: 1

11 0:*  10

11

11 xghost3d 10 Scope: private01  INT

11 Description: Ghostzone size in x direction for 3d arrays

11 Range 10              01  Default: 1

11 0:*  10

11

11 xghost4d 10 Scope: private01  INT

11 Description: Ghostzone size in x direction for 4d arrays

11 Range 10              01  Default: 1

11 0:*  10

11

11 xsize1d 10 Scope: private01  INT

11 Description: Size of 1d array in x direction

11 Range 10              01  Default: 2

11 0:*  10

11

11 xsize2d 10 Scope: private01  INT

11 Description: Size of 2d array in x direction

11 Range 10              01  Default: 2

11 0:*  10

11

11 xsize3d 10 Scope: private01  INT

11 Description: Size of 3d array in x direction

11 Range 10              01  Default: 2

11 0:*  10

11

11 xsize4d 10 Scope: private01  INT

11 Description: Size of 4d array in x direction

11 Range 10              01  Default: 2

11 0:*  10

11 yghost2d 10 Scope: private01  INT

11 Description: Ghostzone size in y direction for 2d arrays

11 Range 10              01  Default: 1

11 0:*  10

11

11 yghost3d 10 Scope: private01  INT

11 Description: Ghostzone size in y direction for 3d arrays

11 Range 10              01  Default: 1

11 0:*  10

11

11 yghost4d 10 Scope: private01  INT

11 Description: Ghostzone size in y direction for 4d arrays

11 Range 10              01  Default: 1

11 0:*  10

11

11 ysize2d 10 Scope: private01  INT

11 Description: Size of 2d array in y direction

11 Range 10              01  Default: 3

11 0:*  10

11

11 ysize3d 10 Scope: private01  INT

11 Description: Size of 3d array in y direction

11 Range 10              01  Default: 3

11 0:*  10

11

11 ysize4d 10 Scope: private01  INT

11 Description: Size of 4d array in y direction

11 Range 10              01  Default: 3

11 0:*  10

11

11 zghost3d 10 Scope: private01  INT

11 Description: Ghostzone size in z direction for 3d arrays

11 Range 10              01  Default: 1

11 0:*  10

11

11 zghost4d 10 Scope: private01  INT

11 Description: Ghostzone size in z direction for 4d arrays

11 Range 10              01  Default: 1

11 0:*  10

11

11 zsize3d 10 Scope: private01  INT

11 Description: Size of 3d array in z direction

11 Range 10              01  Default: 2

11 0:*  10

11 zsize4d 10 Scope: private01  INT

11 Description: Size of 4d array in z direction

11 Range 10              01  Default: 2

11 0:*  10

11

7 Interfaces

General Implements: arrays

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  foo1d  10                    01  compact 11 0
11                    10  gf1d  01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 XGHOST1D
11                    10                    01 group type 11 ARRAY
11                    10                    01  size 11 XSIZE1D
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  foo2d  10                    01  compact 11 0
11                    10  gf2d  01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 XGHOST2D
11 10 x 01 ghostsize 11 YGHOST2D
11                    10                    01 group type 11 ARRAY
11                    10                    01  size 11 XSIZE2D
11 10 x 01 size 11 YSIZE2D
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  foo3d  10                    01  compact 11 0
11                    10  gf3d  01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 XGHOST3D
11 10 x 01 ghostsize 11 YGHOST3D
11                    10                    01  ghostsize 11 ZGHOST3D
11                    10                    01 group type 11 ARRAY
11                    10                    01  size 11 XSIZE3D
11 10 x 01 size 11 YSIZE3D
11                    10                    01  size 11 ZSIZE3D
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  foo4d  10                    01  compact 11 0
11                    10  gf4d  01 dimensions 11 4
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 XGHOST4D
11 10 x 01 ghostsize 11 YGHOST4D
11                    10                    01  ghostsize 11 ZGHOST4D
11                    10                    01  ghostsize 11 TGHOST4D
11                    10                    01 group type 11 ARRAY
11                    10                    01  size 11 XSIZE4D
11 10 x 01 size 11 YSIZE4D
11                    10                    01  size 11 ZSIZE4D
11                    10                    01  size 11 TSIZE4D
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 barxax0d 10 barxax0d 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 barxax1d 10 barxax1d 01 compact 11 0
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1

11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 barxbx1d 10 barxbx1d 01 compact 11 0
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 variable type 11 REAL
11 barxcx1d 10 barxcx1d 01 compact 11 0
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL
11 barxdx1d 10 barxdx1d 01 compact 11 0
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL
11 barxex1d 10 barxex1d 01 compact 11 0
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL
11 barxfx1d 10 barxfx1d 01 compact 11 0
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL
11 barxax2d 10 barxax2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 barxbx2d 10 barxbx2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 variable type 11 REAL
11 barxcx2d 10 barxcx2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 variable type 11 REAL
11 barxdx2d 10 barxdx2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 variable type 11 REAL
11 barxex2d 10 barxex2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL
11 barxfx2d 10 barxfx2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL
11 barxgx2d 10 barxgx2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 barxhx2d 10 barxhx2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL
11 barxix2d 10 barxix2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL
11 barxjx2d 10 barxjx2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL
11 barxkx2d 10 barxkx2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL
11 barxlx2d 10 barxlx2d 01 compact 11 0
11                    10                    01 dimensions 11 2
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL
11 barxax3d 10 barxax3d 01 compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 1
11                    10                    01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 barxbx3d 10 barxbx3d 01 compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 0
11                    10                    01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 variable type 11 REAL
11 barxcx3d 10 barxcx3d 01 compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 1
11                    10                    01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL
11 barxdx3d 10 barxdx3d 01 compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 0
11                    10                    01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL
11 barxex3d 10 barxex3d 01 compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 1
11                    10                    01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL
11 barxfx3d 10 barxfx3d 01 compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 0
11                    10                    01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL
11 barxax4d 10 barxax4d 01 compact 11 0
11                    10                    01 dimensions 11 4
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 1
11                    10                    01  size 11 1
11                    10                    01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 barxbx4d 10 barxbx4d 01 compact 11 0
11                    10                    01 dimensions 11 4
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 0
11                    10                    01  size 11 0
11                    10                    01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 variable type 11 REAL
11 barxcx4d 10 barxcx4d 01 compact 11 0
11                    10                    01 dimensions 11 4
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 1
11                    10                    01  size 11 1
11                    10                    01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL
11 barxdx4d 10 barxdx4d 01 compact 11 0
11                    10                    01 dimensions 11 4
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 0
11                    10                    01  size 11 0
11                    10                    01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 1
11                    10                    01 variable type 11 REAL
11 barxex4d 10 barxex4d 01 compact 11 0
11                    10                    01 dimensions 11 4
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 1
11                    10  x  01  size 11 1
11                    10                    01  size 11 1
11                    10                    01  size 11 1

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL
11 barxfx4d 10 barxfx4d 01 compact 11 0
11                    10                    01 dimensions 11 4
11                    10                    01 distribution 11 DEFAULT
11                    10                    01  ghostsize 11 0
11 10 x 01 ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01  ghostsize 11 0
11                    10                    01 group type 11 ARRAY

11                    10                    01  size 11 0
11                    10  x  01  size 11 0
11                    10                    01  size 11 0
11                    10                    01  size 11 0

11                    10                    01  timelevels 11 1
11                    10                    01 vararrayxsize 11 0
11                    10                    01 variable type 11 REAL

8 Schedule This section lists all the variables which are assigned storage by thorn CactusTest/TestArrays. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 x
11 foo1d foo2d foo3d foo4d  11 x
11 barxax0d  11 x
11 barxax1d barxbx1d barxcx1d barxdx1d barxex1d barxfx1d 11 x
11 barxax2d barxbx2d barxcx2d barxdx2d barxex2d barxfx2d 11 x
11 barxgx2d barxhx2d barxix2d barxjx2d barxkx2d barxlx2d 11 x
11 barxax3d barxbx3d barxcx3d barxdx3d barxex3d barxfx3d 11 x
11 barxax4d barxbx4d barxcx4d barxdx4d barxex4d barxfx4d 11 x

11 x  11 x

11

Scheduled Functions

CCTKxEVOL x usefortranarrays x use fortran arrays

11 x 11 Language: 11 fortran
11 x 11 Sync:  11 barxax0d
11 x 11 x  11 barxax1d
11 x 11 x  11 barxbx1d
11 x 11 x  11 barxcx1d
11 x 11 x  11 barxdx1d
11 x 11 x  11 barxex1d
11 x 11 x  11 barxfx1d
11 x 11 x  11 barxax2d
11 x 11 x  11 barxbx2d
11 x 11 x  11 barxcx2d
11 x 11 x  11 barxdx2d
11 x 11 x  11 barxex2d
11 x 11 x  11 barxfx2d
11 x 11 x  11 barxgx2d
11 x 11 x  11 barxhx2d
11 x 11 x  11 barxix2d
11 x 11 x  11 barxjx2d
11 x 11 x  11 barxkx2d
11 x 11 x  11 barxlx2d
11 x 11 x  11 barxax3d
11 x 11 x  11 barxbx3d
11 x 11 x  11 barxcx3d
11 x 11 x  11 barxdx3d
11 x 11 x  11 barxex3d
11 x 11 x  11 barxfx3d
11 x 11 x  11 barxax4d
11 x 11 x  11 barxbx4d
11 x 11 x  11 barxcx4d
11 x 11 x  11 barxdx4d
11 x 11 x  11 barxex4d
11 x 11 x  11 barxfx4d
11 x 11 Type:  11 function

11

CCTKxINITIAL x setupfortranarrays x set up fortran arrays x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function